Sinalo Bhacela has R200 per week available to spend on beer and cigarettes. Beer costs R10 per bottle and cigarettes cost R40 per packet. Suppose John buys 8 bottles of beer and 3 packets of cigarettes and that his marginal utility from beer is then 20 utils and that from cigarettes 35 utils. Is he in equilibrium? If not, should he buy more beer or more cigarettes? Explain.
1
Expert's answer
2020-11-17T07:19:25-0500

The condition of consumer’s equilibrium is

"\frac{P_{x}}{P_{y}} = \frac{MU_{x}}{MU_{y}}"

In this case

"\frac{10}{40}" is less than "\frac{20}{35}"


  So, John isn’t in equilibrium. He has to buy more beer and less cigarettes, because after that the marginal utility of the last bottle of beer would decrease, while the marginal utility of the last packet of cigarettes would increase, so their ratio would be closer to ¼.